OpenStudy (anonymous):

The mean test score for a class was 70 points, with a standard deviation of 15 points. If I now add 10 points to everyone's test score: a) What is the mean test score for the class after I add the 10 points to each score? b) What is the standard deviation of the test scores for the class after I add the 10 points to each score?

OpenStudy (anonymous):

everything will stay the same because you're not influencing the number of scores that are present, you are just adding 10 to each value of the scores, if you know what i mean

OpenStudy (anonymous):

Kind of sort of. But, we have to show our work on how we can to that conclusion.

OpenStudy (dan815):

ILL SOLVE IT OK

OpenStudy (dan815):

mean changed by 10

OpenStudy (dan815):

std by nothing

OpenStudy (dan815):

since

OpenStudy (dan815):

lets say old mean = sum i from 1 to n of Xi /n

OpenStudy (dan815):

new mean = sum i from 1 to n of Xi+10 /n

OpenStudy (dan815):

= old mean +n*10/n =oldmean+10

OpenStudy (dan815):

so mean has changed by only 10

OpenStudy (dan815):

std = sqrt(sum(diff from mean^2)/n)

OpenStudy (dan815):

the diff from mean didnt change

OpenStudy (dan815):

let k be difference xi-xbar

OpenStudy (dan815):

then xi+10 -(xbar+10) = k still

OpenStudy (dan815):

so you are summing same differences

OpenStudy (anonymous):

That lost me even more.

OpenStudy (dan815):

lol wait

OpenStudy (dan815):

ill draw it after i finish eating

OpenStudy (anonymous):

Ok.

OpenStudy (dan815):